OpenStudy (anonymous):

Express answers in simplest exact form. What is the radius of a circle with a sector area of 7PI sq. ft. and an arc whose measure is 70°? r = ? feet

OpenStudy (anonymous):

What is the formula for the area of a sector?

OpenStudy (anonymous):

Sector = πr^2*degrees/360 7π = πr^2 (70/360 7 = r^2 (7/36) 36 = r^2 so r is either 6 ft or -6 ft

OpenStudy (anonymous):

A negitive radius? That must be beyond the event horizon or something. :)